OpenStudy (amenah8):

HELP!! WILL FAN, MEDAL, AND BEST ANSWER!!!! how would I get y to the opposite side? (y=...)

OpenStudy (amenah8):

\[x=2\sqrt{y+3}\]

OpenStudy (amenah8):

would I square each side????

OpenStudy (anonymous):

yes

OpenStudy (stefrheart):

so first divide both sides by 2

OpenStudy (amenah8):

okay

OpenStudy (anonymous):

is that 2 a specification that it is a square root?

OpenStudy (stefrheart):

\[\frac{ 1 }{ 2 }x=\sqrt{y+3}\]

OpenStudy (amenah8):

no, it is a 2 on the outside of the square root @harmonica1243

OpenStudy (stefrheart):

then you need to square both sides of the equation

OpenStudy (amenah8):

okay

OpenStudy (amenah8):

1/4x=y+3

OpenStudy (stefrheart):

exactly! then you need to subtract the 3 on both sides

OpenStudy (amenah8):

y=1/4x + 3

OpenStudy (stefrheart):

no it should be y=1/4x-3

OpenStudy (amenah8):

oh right, it is subtraction

OpenStudy (stefrheart):

:)

OpenStudy (amenah8):

thank you so much!!!!

OpenStudy (stefrheart):

welcome!!

OpenStudy (stefrheart):

Glad that I helped!